[figure: 9.7 cm]

MAGNETIC PROPERTIES OF SOME RARE-EARTH FERRICYANIDES

R. Göbl^{a}, A. Zentková ^{b}, J. Kováč^{a}, M. Maryško^{c}, A. Zentko^{a} aInstitute of Experimental Physics SAS, Watsonova 47, 043 53 Košice, Slovakia bFaculty of Sciences , P.J. Šafárik University, Košice, Slovakia

cInstitute of Physics, Academy of Sciences of Czech Republic, Prague, Czech Republic

Magnetic properties of rare-earth ferricyanides has been subject of interest for a long time [1]. In this contribution we will present the more detailed information on the structural and magnetic properties of Dy, Sm and Gd ferricyanides. It is found that the compounds investigated are magnetically ordered in the low temperature region (below 15 K). The type of ordering and the value of saturated magnetic moment depends on the type of RE magnetic ion.

Literature

1. F. Hulliger, M. Landolt, H. Vetsch: Journal of Solid State Chemistry 18 (1976) 283

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : Anton Zentko

Address for correspondence : Institute of Experimental Physics SAS Watsonova 47, 043 53 Košice Slovakia

Email address : zentko@saske.sk